OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was four times his age. How old is Sam now? 10 25 20 15

OpenStudy (anonymous):

sam is x years old sam's coach is 2x years old 10 years ago sam was x-10 or y years old sam's coach was 2x-10 or 4y years old

OpenStudy (anonymous):

i dont get how old he is tho

OpenStudy (anonymous):

x-10=y 2x-10=4y how about now?

OpenStudy (anonymous):

for me i start with my choices it could be a bad method but i like it more for these problems

OpenStudy (anonymous):

@MINAMIKE understand that in life, math problem will not always be multiple choice, or multiple choice will not be given, which is why it is better to just understand how to solve it rather than just plugging in the answers and seeing which one works

OpenStudy (anonymous):

also why i hate multiple choice math problems so much, because there is no thinking involved

OpenStudy (anonymous):

you are correct

OpenStudy (anonymous):

if there was no choices i would do it your way

OpenStudy (anonymous):

ineedhelp101 do you still need help

OpenStudy (anonymous):

idk the answer? O.o

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yall wannahelp me with that?

OpenStudy (anonymous):

x-10=y 2x-10=4y 2x-10=4(x-10) solve x

OpenStudy (anonymous):

how did they teach you to solve this problem (system of eqausions x-10=y 2x-10=4y

OpenStudy (anonymous):

-2x=-30?

OpenStudy (anonymous):

and then divide by -2

OpenStudy (anonymous):

you are very close

OpenStudy (anonymous):

15

OpenStudy (anonymous):

yes that is also what i got

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem but completidot did all the work

OpenStudy (anonymous):

any other questions